b.A taxi reaches a police officer while travelling at 30 m/s on the highway. The officer immediately signals the taxi which decelerates uniformly and comes to a complete stop in 5s. Assume zero reaction time.


i.Draw a velocity-time graph to represent the motion of the taxi from the moment the officer

signals to when the taxi comes to a stop. (2)


ii.Determine the distance travelled by the taxi from the time the police signals until the taxi stops. (3)


iii.If the speed limit is 80 km/h, determine whether or not the driver should receive a speeding



1
Expert's answer
2021-02-19T11:50:02-0500

(i)



ii) We have given,


Initial Velocity ,"u= 30m\/s"

time, "t=5s"


Now, using first equation of motion


"v= u+ a\times t"

where, v=final velocity

a=acceleration

Now,

"0=30+a(5)\\\Rightarrow

a=-6m/s^2"


Using second equation of motion


"s=ut+\dfrac{1}{2}(at^2)"


"s=30(5)+\dfrac{1}{2}(-6)(5)^2"


"s=150-75\\

s=75m"


iii) If the speed limit is "30km/h=30\times \dfrac{5}{18}"


"=22.22\text{ms}^{-2}"


Hence, the driver should receive over speeding fine.